x2(d2y/dx2) + x(dy/dx) - 4y = x3

1
Expert's answer
2021-03-18T06:41:43-0400


Given:

"x^2y''+xy'-4y=x^3"


This is an Euler-Cauchy equation with variable coefficients.

Let's try to solve the following homogeneous equation without y-free part (right hand side):

"x^2y''+xy'-4y=\bold{0}"


Try to find any trivial solution in form:

"y=x^a, a - const"

"y'=ax^{a-1}, y'' = a(a-1)x^{a-2}"

"x^2(a(a-1)x^{a-2})+x^2(ax^{a-1})-4x^a=0"


"a(a-1)x^a+ax^a-4x^a=0"

"a^2-a+a-4=0,\space a =\pm2"


We found two trivial solutions, let's choose the first one:

"1. \space y=x^2,\space 2.\space y=x^{-2}"


Now try to find a general solution in form:

"y=x^2u(x)=x^2u"

"y'=2xu+x^2u'"

"y''=2u+4xu'+x^2u''"


"x^2(\blue{2u}+4xu'+x^2u'') + x(\blue{2xu}+x^2u')-\blue{4x^2u}=x^3"

"\blue{4x^3u'}+\underline{x^4}u''+\blue{x^3u'}=x^3"

"5\blue{x^3}u'+\underline{\blue{x^3}x}u''=\blue{x^3}"

"5u'+xu''=1\space(*)"


The equation (*) is the first-order linear differential equation.


Let's solve first homogeneous variant of this equation:

"5u'+xu''=\bold{0}"


To simplify the solution we can use substitution:

"u'(x)=t(x)=t, \space t'=\frac{dt}{dx}=u''"


"5t+xt'=0"

"x\frac{dt}{dx}=-5t"

"\frac{dt}{t}=-5\frac{dx}{x} \space (integrate \space both \space parts)"

"ln|t|=-5ln|x|+ln|C|, \space C - const"

"t=\frac{C}{x^5}"

Now we can use the method of variation of parameters to find the general solution of (*):

"t=\frac{C(x)}{x^5}, \space t'=\frac{C'(x)}{x^5}-5\frac{C(x)}{x^6}"

"\blue{5\frac{C(x)}{x^5}}+x\frac{C'(x)}{x^5}-\blue{5x\frac{C(x)}{x^6}}=1"

"x\frac{C'(x)}{x^5}=1"

"C'(x)=x^4 \space (integrate \space both \space parts)"

"C(x)=\frac{x^5}{5}+D, \space D - const"


"t=\frac{C(x)}{x^5}=\frac{x^5+D}{x^5}=\frac{1}{5}+\frac{D}{x^5}"


"t=u'=\frac{1}{5}+\frac{D}{x^5} \space (integrate \space both \space parts)"

"u=\frac{x}{5}-\frac{D}{4x^4}+E, \space E - const"


"y=x^2u=x^2(\frac{x}{5}-\frac{D}{4x^4}+E)=\frac{x^3}{5}-\frac{D}{4x^2}+Ex^2"

Now we can rewrite constants:

"C_1=-\frac{D}{4}, \space C_2=E"


The general solution of the given equation:

"y=\frac{x^3}{5}+\frac{C_1}{x^2}+C_2x^2, \space C_1,C_2 - const"